answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

From a deck of 52 cards one card is drawn what is the probability that the card will be a red ten?

Since there are 2 red tens in a 52 card deck, probability of drawing one of the 2 red tens is 2/52 or 1/26 or 0.0385.

The probability that a patient recovers from a delicate heart operation is 0.9 what is the probability that exactly 5 of the next 7 patients having the operation survive?

C(7,2)*(.9)^5*(.1)^2, or about .124 = 12.4%

For the desired outcome, considering the seven patients, you need:

(Survive,Survive,Survive,Survive,Survive,Die,Die)

(Survive,Survive,Survive,Survive,Die,Survive,Die)

(Survive,Survive,Survive,Die,Survive,Survive,Die)

.

.

.

(Die,Die,Survive,Survive,Survive,Survive,Survive)

There are C(7,2) [the number of combinations of 7 things taken 2 at a time] = 21 possible desired outcomes. The probability of each of these outcomes is (.9)*(.9)*(.9)*(.9)*(.9)*(.1)*(.1). Multiplying 21 by (.9)*(.9)*(.9)*(.9)*(.9)*(.1)*(.1) yields the answer.

A pair of dice are rolled find the probability at least one dice is a 4?

When rolling one die, the probability of getting a 4 is 1 in 6, or 0.1667. If two dice are rolled, you get two unrelated chances of rolling at least one 4, so the probability is 2 in 6, or 0.3333.

The percentage that is one standard deviation away from mean?

For normally distributed data. One standard deviation (1σ)

Percentage within this confidence interval
68.2689492% (68.3% )


Percentage outside this confidence interval
31.7310508% (31.7% )

Ratio outside this confidence interval
1 / 3.1514871 (1 / 3.15)

What is probability of drawing a seven or a nine out of cards?

Probability of drawing a 7 or 9 is the probability of draw a 7 plus the probability of draw a 9 which is 4/52 + 4/52 = 8/52 = 2/13 = 0.1538.

The fact that probabilities are proportions means that they?

The fact that probabilities are proportions means that they are less than or equal to 1.

What is the probability of a coin landing on its side?

Prob. of Landing on one side = 0.5
Prob. of Landing on the other = 0.5

Therefore, Prob. of Landing on one side or the other = 0.5 +0.5 = 1

:)





I think you mean 'landing on its edge'. In that case it depends on the coin (some even have rounded edges).


Interestingly what face a coin lands on is not exactly 50/50. Each person has an individual flipping-style which leads to a different side being preferred (not explicitly heads vs. tails but in 'landing with the same face upwards as was upwards at the beginning of the flip' vs. 'landing with the other face upwards to the way it started'). On average (with most people) the coin lands the same way up as it started more often than not (you'll have to do hundreds and hundreds of flips before you notice this).




If (like me) you want to be really pedantic you would also take into account the very slight possibility that the coin won'tland at all.

What is the equation of probability?

The probability of an event occurring can be found by dividing the number of favorable outcomes (what you want to happen) by the number of possible outcomes

number of favorable outcomes

probability = _________________________

number of possible outcomes

What is your chance of getting a diamond out of 52 cards?

There is one chance in four of being dealt a diamond from a 52 card deck.

What is the probability of drawing a queen or a club from a deck of cards?

Probability (P) of A or B is: P(A) + P(B) - P(A and B). Apply to the question is: P(Q) + P(Club) - P (Q&Club) = 4/52 + 13/52 - 1/52 = 16/52 = 4/13 or .3077 or 30.77%.

A vertical line is drawn through a normal distribution at z equals plus 0 25 The line separates the distribution into two sections and the smaller section corresponds to of the whole distribution?

At Z = 0.25, the area (larger) to the left of the line is 0.5987. The smaller area to the right of the line is 1 - 0.5987 = 0.4013. Therefore, the smaller section corresponds to 40.13% of the area.

Given a standard normal distriution find the area under the curve which lies to the left of z equals 1.43?

From the table in the related link, Z = 1.43, area to left is 0.9236. This is read directly from the table.

What is the Difference between z scores and percentile ranks?

First let's define both, that will help to see the difference.

1.A percentile is a measure that lets us know what percent of the total frequency scored below that measure. A percentile rank is the percentage of scores that fall below a given score. Here is how that works.

Given a score, call it S and a total of n scores we are looking at, we find the number of scores below S and divide that by n. Next multiply that by 100 and you have the percentile rank.

Now a z score is the number of standard deviations from the mean.
Say the mean is M and your score is S as above. Let sigma be the standard deviation of the distribution. Then z=(S-M)/sigma.

So let's say the mean M is 100 and sigma is 15. S is 132, you did better than average!
So z=(132-100)/15=2.13

If 60 percent of the people scored less than you, then you are in the 60th percentile.
Furthermore, lets say, there were 100 people taking test, then 60 of them scored less than you. Your percentile ranking is (60/100)x100=60

So both are measures of where your results falls in a distribution. z scores are often used for probability of a certain result. Percentile ranks are often used in looking at standardized test results or growth data. One can convert from one to the other.
I have given a conversion table link below

http://www.acposb.on.ca/conversion.htm

How do I find the Z score for tests with mean of 100 and a standard deviation of 20?

Z = (x-mu)/sigma. So, for your example, any x value can be transformed to Z-score by the formula Z = (x-100)/20.

Is binomial distribution is characterized by situations that are analogous to coin tossing?

Yes it is. For a binomial, there must be a fixed number of trials, the probability must remain constant for trials, trials must be independent, and each outcome must be classified into 2 categories.

If you roll two dice what are the odds that you get a one or a five?

I suppose you mean, at least one of those numbers. Just calculate the probability of NOT getting any of those, and take the complement. The probability of not getting a one nor a five on a single die is 4/6 or 2/3. For two dice, the probability is 2/3 x 2/3 = 4/9. So, the probability of getting at least a one or a five with two dice is 1 - 4/9 = 5/9.

I suppose you mean, at least one of those numbers. Just calculate the probability of NOT getting any of those, and take the complement. The probability of not getting a one nor a five on a single die is 4/6 or 2/3. For two dice, the probability is 2/3 x 2/3 = 4/9. So, the probability of getting at least a one or a five with two dice is 1 - 4/9 = 5/9.

I suppose you mean, at least one of those numbers. Just calculate the probability of NOT getting any of those, and take the complement. The probability of not getting a one nor a five on a single die is 4/6 or 2/3. For two dice, the probability is 2/3 x 2/3 = 4/9. So, the probability of getting at least a one or a five with two dice is 1 - 4/9 = 5/9.

I suppose you mean, at least one of those numbers. Just calculate the probability of NOT getting any of those, and take the complement. The probability of not getting a one nor a five on a single die is 4/6 or 2/3. For two dice, the probability is 2/3 x 2/3 = 4/9. So, the probability of getting at least a one or a five with two dice is 1 - 4/9 = 5/9.

What is 2 thirds times 4?

In 4/3 * 2, just multiply th numerator by 2, to get 8/3.

If you role 2 fair dice and add the scores which is the most likely number?

The most likely number is 7 since and here are the ways to get it:

1+6, 2+5, 3+4, 4+3, 5+2, 6+1.

Trending Questions
When did Samantha Rolls die? What is the probability of getting a 0 after rolling a single die numbered 1 to 6? How many cards are in a five crown deck? Does smoking weed produce more girl babies than boys? How are sampling distributions generated using the empirical sampling approach? If a die is rolled 6 times what is the probability of 3 appearing exactly 2 times? What does it mean to get prosecuted? Why the probability of getting hurt is more when a man jumps from significant heidht? If a normal woman whose father is color-blind marries a colorblind man what are their chances of having a colorblind son? How many combinations of two numbers are there in 16 numbers? How many ways can you choose 5 hearts from a 52-card deck? What is the probability of choosing a higher card than the 6 of diamonds? Which Die - Coin combinations' experimental probability matches its theoretical probability? What is the probability of not getting a 2 or not getting a 5? Bogart's classics marbles there are 18 in a bag 1 large 17 small ones never been open whats there worth? How many outcomes are possible when two dice are rolled.? How many 4 digit EVEN numbers have all 4 digits distinct? How many possible hands are there for getting 3 Cards of the same suit and 2 random cards? What is the probability of a die landing with a 3 up? A box contains three coins 2 of them fair and one two headed coin select at random tossed twice heads appear both time what is the probability that the coin is two headed?